# Complements of discriminants of real parabolic function singularities. II

### 08 Mar 2026 v5

@@ Abstract.
We list all local connected components of sets of non-discriminant functions near all parabolic function singularities (which are the second most important family of singularity classes of smooth functions after simple singularities). Thus, we prove (and improve in one particular case) all the corresponding conjectures from the previous work [25] with the same title. As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs, in particular, find a new one. We also show that the complements of the discriminant varieties of the $X_{9}^{\pm}$ and $P_{8}^{1}$ singularities within their versal deformations have nontrivial one-dimensional homology groups, in contrast to all simple singularities.
These results are applications of a general method for investigating and separating non-singular perturbations of real function singularities. An important part of this method is a computer program that formalizes Picard–Lefschetz theory and surgeries of Morse functions.
@@
discriminant, singularity, surgery, versal deformation, isotopy classification
@@ 1.1. Motivation and terminology
Many functions studied in these theories are defined by integral transforms. Their arguments are the parameters of certain function families (called generating families), and their values are defined by the integrals of certain differential forms along the cycles in the zero-level sets of the corresponding functions. These integral functions are regular outside the discriminant varieties of the families. The analytic behavior of such a function near a discriminant point depends on the singularity type of the corresponding function and on the connected component of the complement of the discriminant variety, from which we approach this point. Thus, enumerating these components is one of the first steps in studying the analytical properties of the integral transforms.
@@ 1.2. Structure and interpretation of results
Each class of parabolic singularities is unimodal, that is, it is swept out by a one-parametric family of orbits of the natural action of the group of local diffeomorphisms $({\mathbb{R}}^{n},0)\to({\mathbb{R}}^{n},0)$. In each case, we consider the canonical versal deformation of the corresponding class. This is a space of polynomials that is it diffeomorphic to ${\mathbb{R}}^{9}$ (respectively, ${\mathbb{R}}^{10}$, ${\mathbb{R}}^{8}$) in the case of singularities of classes $X_{9}^{*}$, ($J_{10}^{*}$, $P_{8}^{*}$), see formulas (1), (31), (32), (42). The polynomials with parabolic singularities of the corresponding class form a subset that is diffeomorphic to an open interval in this space. All orbits are represented by some points of these sets. Our main results are the lists of connected components, into which the discriminant varieties divide these spaces, see Theorems 6, 14, 19, 29, 32, 39, and 43. However, each of these results can be interpreted as a list of all local components of the discriminant complements in arbitrarily small neighborhoods of an arbitrary particular parabolic singularity of the corresponding class. Indeed, one-parametric groups of homogeneous or weighted homogeneous dilations act on the parameter spaces of versal deformations. These groups are naturally isomorphic to the multiplicative group ${\mathbb{R}}^{1}_{+}$ of positive numbers. For $P_{8}$ singularities the element $\{t\}\in{\mathbb{R}}^{1}_{+}$ maps each of degree three polynomial $f_{\lambda}(x,y,z)$ to $t^{-3}f_{\lambda}(tx,ty,tz)$, in particular preserves the principal homogeneous part of the polynomials. For $X_{9}$ singularities it acts as $\{t\}:f_{\lambda}(x,y)\to t^{-4}f_{\lambda}(tx,ty)$, and for $J_{10}$ as $\{t\}:f_{\lambda}(x,y)\to t^{-6}f_{\lambda}(t^{2}x,ty)$. All of these actions preserve the discriminants and their complements. Parabolic singularities are the only invariant points of these actions. Each point of the discriminant complement is reduced by this action into an arbitrarily small neighborhood of some parabolic singularity. Additionally, according to [17], the discriminant variety is equisingular along the one-dimensional sets of parabolic singularities in parameter spaces of versal deformations. Therefore, each component of the discriminant complement, which is represented in an arbitrarily small neighborhood of some particular parabolic singularity, is also represented in the neighborhoods of all other parabolic singularities in the one-parametric sets.
@@ 1.3. On the method
In [25], an invariant of the connected components of the complements of the discriminant varieties of function singularities is described, complete lists of possible values of this invariant for parabolic singularities are found, and all of these values are realized by concrete polynomials. The remaining problem is to list all components with each value. In [25], a conjectural answer to this problem was proposed. It provides the lists consisting of 7, 15, 7, 14, 52, 11, and 33 connected components for the singularities of classes $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively. Below, we prove that almost all of these lists are indeed complete. However, some two sets of perturbations of the $J_{10}^{1}$ singularities indicated in [25] consist of pairs of different connected components. Thus, the number 11 is replaced in Table 2 by 13.
@@ 1.4. On the discriminant complements for simple singularities
E. Looijenga [16] proved that all local connected components of discriminant complements of real simple singularities are homotopically trivial. Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 13 and 41.
@@
In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [2], [9], [14], [23]). Near the non-singular points of wavefronts (corresponding to the Morse singularities of generating functions) the local lacunas were enumerated in [7] and [6], see also [2]. Near the simplest non-Morse points of classes $A_{2}$ and $A_{3}$, this was done in [9], and near all simple singularities in [21]. In [24], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that this list is complete. We prove this conjecture for almost all parabolic classes but discover one additional local lacuna in the $P_{8}^{2}$ case, see § 5.
@@ 2.2. Virtual functions and virtual components
A virtual function associated with a generic real polynomial $f$ is a collection of topological characteristics of the complex zero-level set of this polynomial. Specifically, the data of a generic polynomial $f(x,y)$ of some class $X_{9}^{*}$ consists of
@@
    the $9\times 9$ matrix of the intersection indices of canonically ordered and oriented vanishing cycles forming the basis of the two-dimensional homology group of the submanifold of ${\mathbb{C}}^{3}$ given by the equation $f(x,y)+z^{2}=0$,
@@
    the string of intersection indices of these nine cycles with the set of real points,
@@
In the statement of Proposition 2 of [28], the functions $f$ and $\tilde{f}$ belong to the broader class of all degree-four Morse polynomials with non-degenerate principal homogeneous parts. However, the entire construction proving this proposition in [28] can be performed within the spaces of the versal deformations (1). Thus, it also proves the above Proposition 3.
@@ 2.3. Enumeration of components for $X_{9}^{+}$ singularities
There are exactly seven connected components in the space of non-discriminant functions of class $\Phi_{+}$. The shapes of the zero-level sets of some polynomials representing these components are shown in Fig. 1.
@@
Using the Lyashko–Looijenga covering (in the form described in [12]), we can assume that the polynomials $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the common associated virtual function. Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has in its interior points only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the union of the discriminant and the set of polynomials with fewer than nine critical values). Since the virtual function of $\tilde{f}$ is the same as that of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$. (For this existence statement, the parabolicity of $X_{9}$ singularities is essential. The proof of it is based on Proposition 2 of [12].) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2). First, suppose that it is the same point $I(1)$. A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that preserves the associated virtual functions and commutes with the Lyashko–Looijenga map. This automorphism is induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$, see [28]. Therefore, there are only two possibilities: either the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are either the same or mapped to each other by this automorphism. In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this automorphism, which belongs to the group $G_{0}$. If the point $\tilde{I}(1)$ is different from $I(1)$, we first apply the element of the group $G_{0}$, moving these points to each other, to the entire path $\tilde{I}$ (in particular, to its starting point $\tilde{f}$). This leads us to the situation considered previously. $\Box$
@@ 2.4. The 1-cohomology of discriminant complements of $X_{9}^{+}$ singularities
Recall the notation $B({\mathbb{R}}^{2},k)$ for the unordered configuration space, whose points are arbitrary $k$-element subsets of the plane. For any $k\geq 2$, $H^{1}(B({\mathbb{R}}^{2},k))\simeq{\mathbb{Z}}$. The basic element of this group is induced from $H^{1}({\mathbb{C}}^{1}\setminus\{0\})$ by the map $B({\mathbb{R}}^{2},k)\equiv B({\mathbb{C}}^{1},k)\to{\mathbb{C}}^{1}\setminus\{0\}$ sending any collection of $k$ distinct point $z_{1},\dots,z_{k}$ to the product of all k(k1)k(k-1) complex numbers $z_{i}-z_{j}$, $i\neq j$. Let $U$ be a connected component of the discriminant complement in $\Phi_{+}$ that consists of polynomials whose zero-level sets are unions of $k\geq 2$ non-nested ovals. Then a map $H^{*}(B({\mathbb{R}}^{2},k))\to H^{*}(U)$ is defined as follows. For any polynomial $f\in U,$ define $\beta(f)$ as the space of all choices of a point inside any of its ovals. These spaces $\beta(f)$ form a locally trivial fiber bundle over the component $U$. The fibers of this fiber bundle are contractible, hence it has a cross-section, which is a continuous map $U\to B({\mathbb{R}}^{2},k)$. Clearly, these maps defined by different cross-sections are homotopic and define the same homomorphism of cohomology rings.
@@ 2.5. The components for the $X_{9}^{1}$ singularities
acts on the space $\Phi_{1}$. It preserves the discriminant variety and hence permutes the connected components of its complement. This involution naturally extends to the set of associated virtual components: each virtual component which is the value of the set-valued invariant of the polynomial $f(x,y)$ goes to the value of the polynomial f(x,y)-f(x,-y). This involution has no fixed elements: indeed, the sum of the values of the invariant Ind (see Definition 3) for any two components (real or virtual) related by the involution is always equal to 1-1.
@@
Proof. Let $f$ be a polynomial that satisfies the conditions of this lemma. By Lemma 5 of [28], the $j$-invariant of its principal homogeneous part is equal to 0, hence $A=\pm\frac{\sqrt{3}}{2}$ in (12). We can assume that the coordinates $(x,y)$ are centered at the $E_{7}$ point. In these coordinates, the monomials of degree $\leq 2$ of the polynomial $f$ vanish, and the homogeneous part of degree three is proportional to the third degree of a linear function that also divides the degree four homogeneous part of $f$. Thus, the degree three homogeneous part has the form $\alpha x^{3}$ or $\alpha y^{3}$, $\alpha\neq 0$. The critical value of the polynomial at the $A_{2}$ point is equal to $\pm 3^{\frac{15}{2}}2^{-8}\alpha^{2}$, where $\pm$ is the sign of $A$. This value is equal to $3^{\frac{15}{2}}2^{-8}$ if and only if this sign is ++ and $\alpha$ equals $1$ or 1-1. $\Box$
Proof of Proposition 16. The deduction of this proof from Lemma 17 essentially repeats the proof of Proposition 9 with the following simplification. In this case, the neighborhood of the point ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}+x^{3})$ in the parameter space of the deformation (1) has no nontrivial automorphisms that commute with the Lyashko–Looijenga covering (see [28], § 4.5). Therefore, the statement of Proposition 16 follows immediately from the fact that all four polynomials (14) belong to the same orbit of the group $G_{1}$. $\Box$
@@
$xy(x^{2}+y^{2}-t)-\alpha t^{2},\quad\alpha\in(0,1/8)$

### 16 Mar 2026 v6

@@ Abstract.
We list all connected components of sets of non-discriminant functions near all parabolic function singularities (which are the second most important family of singularity classes of smooth functions after simple singularities). Thus, we prove (and improve in one particular case) all the corresponding conjectures from the previous work [26] with the same title. As an application, we enumerate all local Petrovskii lacunas near arbitrary parabolic singularities of wavefronts of hyperbolic PDEs. We also show that the complements of the discriminant varieties of the versal deformations of $X_{9}^{\pm}$ and $P_{8}^{1}$ singularities have nontrivial one-dimensional homology groups, in contrast to all simple singularities.
These results are applications of a general method for investigating and separating non-singular perturbations of real function singularities. An important part of this method is a computer program that formalizes local Picard–Lefschetz theory and surgeries of Morse functions.
@@
discriminant, singularity, surgery, versal deformation, isotopy classification, Lyashko–Looijenga covering, Picard–Lefschetz theory
@@ 1.1. Motivation and terminology
Many functions studied in these theories are defined by integral transforms. Their arguments are the parameters of certain function families (called generating families), and their values are defined by the integrals of certain differential forms along the cycles in the zero-level sets of the corresponding functions. These integral functions are regular outside the discriminant varieties of the generating families. The analytic behavior of such a function near a discriminant point depends on the singularity type of the corresponding function and on the connected component of the complement of the discriminant variety, from which we approach this point. Thus, enumerating these components for different types of discriminant points is one of the first steps in studying the analytical properties of the integral transforms.
@@ 1.2. Structure and interpretation of results
Each class of parabolic singularities is unimodal, that is, it is swept out by a one-parametric family of orbits of the natural action of the group of local diffeomorphisms $({\mathbb{R}}^{n},0)\to({\mathbb{R}}^{n},0)$. In each case, we consider the canonical versal deformation of the corresponding class. This is a space of polynomials that is it diffeomorphic to ${\mathbb{R}}^{9}$ (respectively, ${\mathbb{R}}^{10}$, ${\mathbb{R}}^{8}$) in the case of singularities of classes $X_{9}^{*}$, (respectively, $J_{10}^{*}$, $P_{8}^{*}$), see formulas (1), (31), (32), (42). The polynomials with parabolic singularities of the corresponding class form a subset that is diffeomorphic to an open interval in this space. All orbits are represented by some points of these intervals. Our main results are the lists of connected components, into which the discriminant varieties divide the parameter spaces of these deformations, see Theorems 6, 14, 19, 30, 33, 40, and 44. However, each of these results can be interpreted as a list of all local components of the discriminant complements in arbitrarily small neighborhoods of an arbitrary particular parabolic singularity of the corresponding class. Indeed, one-parametric groups of homogeneous or weighted homogeneous dilations act on the parameter spaces of canonical versal deformations. These groups are naturally isomorphic to the multiplicative group ${\mathbb{R}}^{1}_{+}$ of positive numbers. For $P_{8}$ singularities the element $\{t\}\in{\mathbb{R}}^{1}_{+}$ maps each of degree three polynomials $f_{\lambda}(x,y,z)$ to $t^{-3}f_{\lambda}(tx,ty,tz)$, in particular it preserves the principal homogeneous part of the polynomials. For $X_{9}$ singularities it acts as $f_{\lambda}(x,y)\to t^{-4}f_{\lambda}(tx,ty)$, and for $J_{10}$ as $f_{\lambda}(x,y)\to t^{-6}f_{\lambda}(t^{2}x,ty)$. All of these actions preserve the discriminants and their complements. Parabolic singularities are defined by the homogeneous or weighted homogeneous polynomials of top degree and are the only invariant points of these actions. Each point of the discriminant complement is reduced by this action into an arbitrarily small neighborhood of some parabolic singularity. Additionally, according to [17], the discriminant variety is equisingular along the one-dimensional sets of parabolic singularities in parameter spaces of versal deformations. Therefore, each component of the discriminant complement, which is represented in an arbitrarily small neighborhood of some particular parabolic singularity, is also represented in the neighborhoods of all other parabolic singularities of the same class.
@@ 1.3. On the method
In [26], an invariant of the connected components of the complements of the discriminant varieties of function singularities is described, complete lists of possible values of this invariant for all parabolic singularities are found, and all of these values for singularities of classes $X_{9}$ and $J_{10}$ are realized by concrete polynomials. The remaining problems are to list all components with each value of the invariant, and also to realize them for the $P_{8}$ singularities. In [26], a conjectural answer to this problem was proposed. It provides the lists consisting of 7, 15, 7, 14, 52, 11, and 33 connected components for the singularities of classes $P_{8}^{1}$, $P_{8}^{2}$, $X_{9}^{\pm}$, $X_{9}^{1}$, $X_{9}^{2}$, $J_{10}^{1}$, and $J_{10}^{3}$ respectively. Below, we prove that almost all of these lists are indeed complete. However, some two sets of perturbations of the $J_{10}^{1}$ singularities indicated in [26] consist of pairs of different connected components. Thus, the number 11 is replaced in Table 2 by 13.
@@ 1.4. On the discriminant complements for simple singularities
E. Looijenga [16] proved that all local connected components of discriminant complements of real simple singularities are homotopically trivial. Conversely, we demonstrate that at least three such components of parabolic singularities of each of the classes $X_{9}^{+}$ and $X_{9}^{-}$, as well as one component of the $P_{8}^{1}$ class, have nontrivial one-dimensional homology groups; see Theorems 13 and 42.
@@
In the theory of hyperbolic PDEs, local Petrovskii lacunas are the local connected components of wavefront complements where wave functions behave regularly (see, for example, [2], [9], [14], [19], [24]). Near the non-singular points of wavefronts (corresponding to the Morse singularities of generating functions) the local lacunas were enumerated in [7] and [6], see also [2]. Near the simplest non-Morse points of classes $A_{2}$ and $A_{3}$, this was done in [9], and near all simple singularities in [22]. In [25], a list of Petrovskii lacunas at parabolic singular points of wavefronts was presented, and it was conjectured that this list is complete. We prove this conjecture for almost all parabolic classes but discover one additional local lacuna in the $P_{8}^{2}$ case, see § 5.
@@ 2.2. Virtual functions and virtual components
A virtual function associated with a generic real polynomial $f$ is a collection of topological characteristics of the complex zero-level set of this polynomial. Specifically, these data of a generic polynomial $f(x,y)$ of some class $X_{9}^{*}$ consists of
@@
    the $9\times 9$ matrix of the intersection indices of canonically ordered and oriented vanishing cycles forming the basis of the two-dimensional homology group of the submanifold $V_{f}\subset{\mathbb{C}}^{3}$ given by the equation $f(x,y)+z^{2}=0$,
@@
    the string of intersection indices of these nine cycles with the set of real points in $V_{f}$,
@@
In the statement of Proposition 2 of [29], the functions $f$ and $\tilde{f}$ belong to the broader class of all degree-four Morse polynomials with non-degenerate principal homogeneous parts. However, the entire construction proving this proposition in [29] is performed within the spaces of the versal deformations (1). Thus, it also proves the above Proposition 3.
@@ 2.3. Enumeration of components for $X_{9}^{+}$ singularities
There are exactly seven connected components in the space of non-discriminant polynomials of class $\Phi_{+}$. The shapes of the zero-level sets of some polynomials representing these components are shown in Fig. 1.
@@
Using the Lyashko–Looijenga covering (in the form described in [12]), we can assume that the polynomials $f$ and $\tilde{f}$ have the same set of critical values and the same set of paths in ${\mathbb{C}}^{1}$ specifying the common associated virtual function. Let $I:[0,1]\to\Phi_{+}$ be a generic path such that $I(0)=f$, $I(1)$ is the polynomial ${\bf T}(x^{4}+x^{3}+y^{4})$, and the path $I$ has in its interior points only finitely many transversal intersections with the variety of non-generic polynomials in $\Phi_{+}$ (that is, with the union of the discriminant and the set of polynomials with fewer than nine critical values). Since the virtual function of $\tilde{f}$ is the same as that of $f$, there exists another path $\tilde{I}:[0,1]\to\Phi_{+}$ such that $\tilde{I}(0)=\tilde{f}$, the polynomials $I(t)$ and $\tilde{I}(t)$ have the same set of critical values for any $t\in[0,1]$, and the polynomials of the path $\tilde{I}$ undergo the surgeries of the same type as $I$ at the same instants $t$. (For this existence statement, the parabolicity of $X_{9}$ singularities is essential. The proof of it is based on Proposition 2 of [12].) In particular, the point $\tilde{I}(1)$ can only be one of the four points (2). First, suppose that it is the same point $I(1)$. A neighborhood of this point in the space $\Phi_{+}$ has only one nontrivial smooth automorphism that preserves the associated virtual functions and commutes with the Lyashko–Looijenga map. This automorphism is induced by the reflection of ${\mathbb{R}}^{2}$ in the line $y=0$, see [29]. Therefore, there are only two possibilities: the final segments of the paths $I$ and $\tilde{I}$ (for $t\in(1-\varepsilon,1]$) are either the same or are mapped to each other by this automorphism. In both cases, this condition can be continued to the entire paths $I$ and $\tilde{I}$ and proves that either $\tilde{f}=f$ or these two functions are mapped to each other by this automorphism, which belongs to the group $G_{0}$. If the point $\tilde{I}(1)$ is different from $I(1)$, we first apply the element of the group $G_{0}$, moving these points to each other, to the entire path $\tilde{I}$ (in particular, to its starting point $\tilde{f}$). This leads us to the situation considered previously. $\Box$
@@ 2.4. The 1-cohomology of discriminant complements of $X_{9}^{+}$ singularities
Recall the notation $B({\mathbb{R}}^{2},k)$ for the unordered configuration space, whose points are arbitrary $k$-element subsets of the plane. For any $k\geq 2$, $H^{1}(B({\mathbb{R}}^{2},k))\simeq{\mathbb{Z}}$. The basic element of this group is induced from the generator of the group $H^{1}({\mathbb{C}}^{1}\setminus\{0\})\sim{\mathbb{Z}}$ by the map $B({\mathbb{R}}^{2},k)\equiv B({\mathbb{C}}^{1},k)\to{\mathbb{C}}^{1}\setminus\{0\}$ sending any collection of $k$ distinct point $z_{1},\dots,z_{k}$ to the product of all k(k1)k(k-1) complex numbers $z_{i}-z_{j}$, $i\neq j$. Let $U$ be a connected component of the discriminant complement in $\Phi_{+}$ that consists of polynomials whose zero-level sets are unions of $k\geq 2$ non-nested ovals. Then a map $H^{*}(B({\mathbb{R}}^{2},k))\to H^{*}(U)$ is defined as follows. For any polynomial $f\in U,$ define $\beta(f)$ as the space of all choices of a point inside any of its ovals. These spaces $\beta(f)$ form a locally trivial fiber bundle over the component $U$. The fibers of this fiber bundle are contractible, hence it has a cross-section, which is a continuous map $U\to B({\mathbb{R}}^{2},k)$. Clearly, these maps defined by different cross-sections are homotopic and define the same homomorphism of cohomology rings.
@@ 2.5. The components for the $X_{9}^{1}$ singularities
acts on the space $\Phi_{1}$. It preserves the discriminant variety and permutes the connected components of its complement. This involution naturally extends to the set of associated virtual components: each virtual component which is the value of the set-valued invariant of the polynomial $f(x,y)$ goes to the value of the polynomial f(x,y)-f(x,-y). This involution has no fixed elements: indeed, the sum of the values of the invariant Ind (see Definition 3) for any two components (real or virtual) related by the involution is always equal to 1-1.
@@
Proof. Let $f$ be a polynomial that satisfies the conditions of this lemma. By Lemma 5 of [29], the $j$-invariant of its principal homogeneous part is equal to 0. By the formula for $j$-invariants of quartic forms (see [18], §1.3), $A=\pm\frac{\sqrt{3}}{2}$ in (12). We can assume that the coordinates $(x,y)$ are centered at the $E_{7}$ point. In these coordinates, the monomials of degree $\leq 2$ of the polynomial $f$ vanish, and the homogeneous part of degree three is proportional to the third degree of a linear function that also divides the degree four homogeneous part of $f$. Thus, the degree three homogeneous part has the form $\alpha x^{3}$ or $\alpha y^{3}$, $\alpha\neq 0$. The critical value of the polynomial at the $A_{2}$ point is equal to $\pm 3^{\frac{15}{2}}2^{-8}\alpha^{2}$, where $\pm$ is the sign of $A$. This value is equal to $3^{\frac{15}{2}}2^{-8}$ if and only if this sign is ++ and $\alpha$ equals $1$ or 1-1. $\Box$
Proof of Proposition 16. The deduction of this proof from Lemma 17 essentially repeats the proof of Proposition 9 with the following simplification. In this case, the neighborhoods of the point ${\bf\tilde{T}}(x^{3}y+\sqrt{3}x^{2}y^{2}+xy^{3}+x^{3})$ in the parameter space of the deformation (1) have no nontrivial automorphisms that commute with the Lyashko–Looijenga covering (see [29], § 4.5). Therefore, the statement of Proposition 16 follows immediately from the fact that all four polynomials (14) belong to the same orbit of the group $G_{1}$. $\Box$
@@
$xy(x^{2}+y^{2}-t)-\alpha t^{2},\quad\alpha\in(0,1/8),$